Ask your own question, for FREE!
Algebra 6 Online
OpenStudy (angel_kitty12):

PLEASE HELP ASAP (Algebra 2 Conditional Probability) Thanks! Draw a tree diagram. Find the probability. A softball game has an 80% chance of being canceled for a light drizzle and a 30% chance of being canceled for a heavy fog when there is no drizzle. There is a 70% chance of heavy fog and a 30% chance of light drizzle. A. Find the probability that the game will be canceled. B. Find the probability there will be a light drizzle and the game will not be canceled. PLEASE HELP I DONT UNDERSTAND WHAT I MUST DO TO FIND THE ANSWERS!

OpenStudy (kropot72):

|dw:1445137222005:dw| From the probability tree drawn above: The probability of the game being canceled given a heavy fog is 0.7 * 0.3 = 0.21. The probability of the game being canceled given a light drizzle is 0.3 * 0.8 = 0.24. A. The probability of the game being canceled is the sum of the above two values of conditional probability: 0.21 + 0.24 = 0.45.

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
15 hours ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
18 hours ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Arizona: does anyone know the percentage rate of 30,000 and 40,000
1 week ago 0 Replies 0 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!